OpenStudy (anonymous):

y varies inversely as twice x When x = 4, y = 1. Find y when x = 5.

OpenStudy (mertsj):

\[y=\frac{k}{2x}\]

OpenStudy (mertsj):

Replace x with 4 and y with 1 and find k.

OpenStudy (mertsj):

What did you get for k?

OpenStudy (anonymous):

k = 2x * y = 2 *4 * 1 = 8 If x = 5, y = k/2x = 8/ 2* 5 = .8